# On the Convergence Rate of LoRA Gradient Descent

### v1

@@ Abstract
The low-rank adaptation (LoRA) algorithm for fine-tuning large models has grown popular in recent years due to its remarkable performance and low computational requirements. LoRA trains two “adapter” matrices that form a low-rank representation of the model parameters, thereby massively reducing the number of parameters that need to be updated at every step. Although LoRA is simple, its convergence is poorly understood due to the lack of Lipschitz smoothness, a key condition for classic convergence analyses. As a result, current theoretical results only consider asymptotic behavior or assume strong boundedness conditions which artificially enforce Lipschitz smoothness. In this work, we provide for the first time a non-asymptotic convergence analysis of the original LoRA gradient descent algorithm, which reflects widespread practice, without such assumptions. Our work relies on three key steps: i) reformulating the problem in terms of the outer product of the stacked adapter matrices, ii) a modified descent lemma for the “Lipschitz-like” reparametrized function, and iii) controlling the step size. With this approach, we prove that LoRA gradient descent converges to a stationary point at rate $O(\frac{1}{\log T})$, where $T$ is the number of iterations.
@@ 1 Introduction
Modern applications of large language models (LLMs) typically involve self-supervised pretraining of a large foundation model, which is later fine-tuned on a smaller task-specific dataset through supervised learning. Parameter efficient fine-tuning methods, which only train a small number of additional parameters, can efficiently adapt LLMs to a multitude of downstream tasks while maintaining comparable performance to fine-tuning of the full model parameters. One highly popular method is LoRA, or Low-Rank Adaptation [hu2022lora], which trains a low-rank representation of the model parameters. Specifically, once the pretraining stage yields a model weight matrix $W_{0}$, LoRA only updates low-rank adapter matrices $A$, $B$, such that the final weights are $W_{0}+BA$. The matrices $A$ and $B$ can be updated with any optimization method, including gradient descent. However, even this simple algorithm is challenging to analyze; as observed in [sun2024improving] and [malinovsky2024randomizedasymmetricchainlora], even if the original loss function is Lipschitz smooth, the new loss function— reparametrized in terms of $A$ and BB— is not. This nonsmoothness is a major obstacle to applying classic convergence proof techniques of gradient methods to LoRA.
As a result, current theoretical analyses of LoRA fall into one of three categories. First, there are several works that examine LoRA in infinite regimes, such as its asymptotic convergence or its behavior on infinitely wide neural networks [kim2025lora, jang2024lora, zeng2024the, yaras2024compressible, zhang2024riemannianpreconditionedlorafinetuning, hayou2024loraefficientlowrank]. However, these works do not establish a non-asymptotic convergence rate for finite models. Second, many works propose and analyze memory-efficient algorithms that resemble but do not address the original LoRA algorithm. These include GaLore [zhao2024galore] and LoRA variants that only update a single adapter matrix at a time [malinovsky2024randomizedasymmetricchainlora, sokolov2025bernoulliloratheoreticalframeworkrandomized]. Third, some recent works analyze algorithmic frameworks that do extend to LoRA, and do derive a convergence rate [jiang2024a, sun2024improving, ghiasvand2025decentralizedlowrankfinetuninglarge]. However, these works require a strong boundedness assumption that artificially enforces Lipschitz smoothness within the new parametrization. Typically, they require that the adapter matrices $A$ and $B$ are uniformly upper bounded by some constants which appear in the final convergence bound. Considering the existing literature, we seek to answer the question:
@@
In this work, we close the above research gap by establishing the first non-asymptotic convergence analysis of the LoRA gradient descent algorithm, without requiring asynchronous updates, bounded adapter matrices, or Lipschitz smoothness of the reparametrized loss function. We only assume that the original loss function is Lipschitz smooth. Our proof approach involves stacking $B$ and $A$ into a single matrix $V$ and deriving a modified descent lemma for gradient descent with respect to $V$. We show that to achieve descent at each step, the algorithm requires a step size that is “small enough” with respect to the norm of the current parameters and gradient. This recursive relationship between the parameters, gradient, and step size introduces instability in the convergence, resulting in an $O(\frac{1}{\log T})$ convergence rate to a stationary point. If we further assume that the adapter norms are bounded, we can recover the classic $O(\frac{1}{T})$ convergence rate of gradient descent on Lipschitz smooth functions.
Our theoretical analysis reveals two key insights. First, the convergence rate of LoRA does not (and should not) depend on the chosen rank, since gradient descent is a dimension-free algorithm. Second, without additional structural assumptions, the point of convergence of LoRA can be arbitrarily far away from that of full-rank gradient descent. This supports empirical observations that LoRA can yield completely different solutions than full-rank finetuning [shuttleworth2025loravsfinetuningillusion, biderman2024lora].
In Section 2, we review the related work in detail and contextualize our contribution. In Section 3, we provide our convergence analysis, establishing preliminaries in Section 3.1 and providing our main results in Section 3.2. Finally, in Section 4, we discuss the insights from our work.
@@ 2 Related Work
The LoRA algorithm, first proposed in [hu2022lora], is motivated by the idea of a viable “low-rank representation” of deep neural networks [oymaklowrank, li2018measuring]. In particular, [aghajanyan-etal-2021-intrinsic] argues that large overparametrized models reside on a lower “intrinsic dimension,” and fine-tuning with a low dimension reparametrization, such as a random subset of the weight matrices, can yield strong performance while minimizing computational expenses.
@@
LoRA is simple, popular, and empirically effective for finetuning large language models. Many variants have been proposed to improve the initialization, performance, memory efficiency, or privacy of the algorithm [shen2025kroneckerlorahybridkroneckerloraadapters, büyükakyüz2024oloraorthonormallowrankadaptation, li2025flatlora, meng2024pissa, wang2024loraga], including but not limited to federated LoRA [10.24963/ijcai.2025/1196], quantized LoRA (QLoRA) [dettmersqlora], and ReLoRA [lialin2024relora]. While LoRA performs well in practice, its behavior in comparison to full-rank training is poorly understood, with empirical studies showing that the two approaches can produce completely different solutions [shuttleworth2025loravsfinetuningillusion, biderman2024lora]. Consequently, the theoretic properties of LoRA are highly relevant for understanding the strengths and weakness of this method.
@@
As highlighted in the introduction, the theory of LoRA convergence has been tackled from many angles. First, many works examine the behavior of LoRA in infinite regimes, whether by characterizing the kinds of solutions LoRA converges to at infinity or analyzing its convergence in the neural tangent kernel (NTK) regime, which defines the linearized training dynamics of an infinitely wide neural network [pmlr-v202-malladi23alazy]. For example, [kim2025lora] argues that LoRA converges to a global minimum with high probability, as long as such a low-rank minimizer exists. The work [zeng2024the] studies the expressive power of the low-rank solutions achieved by LoRA. The work [jang2024lora] shows that LoRA has no spurious local minima in the NTK regime. The LoRA+ [hayou2024loraefficientlowrank] algorithm, which updates the adapter matrices with differently scaled learning rates, is motivated by an NTK convergence analysis. Finally, the gradient flow dynamics of LoRA have been studied specifically for matrix factorization [xu2025understandingGF]. These works do not establish a convergence rate for discrete-time LoRA for general functions.
@@
Second, there are memory-efficient LoRA-like algorithms with concrete convergence analyses that do not apply to the original LoRA formulation. These include GaLore, an algorithm which projects the gradient into a low-rank subspace but maintains a full-rank update of the parameters [zhao2024galore], LDAdam [robert2025ldadam], which performs adaptive optimization steps within lower dimensional subspaces, and Randomized Subspace Optimization (RSO) [chen2025arso]. The LoRA variant LoRA-One incorporates information from the full gradient at initialization [zhang2025loraone], and it is analyzed only for Gaussian data and mean-squared loss. Finally, the variants RAC-LoRA [malinovsky2024randomizedasymmetricchainlora] and Bernoulli-LoRA [sokolov2025bernoulliloratheoreticalframeworkrandomized] only train one adapter matrix while freezing the other as a fixed projection matrix, essentially preserving Lipschitz smoothness while failing to capture the complex training dynamics of actual LoRA implementations. In practice, the LoRA matrices are updated simultaneously, causing nonlinearity and nonsmoothness in the optimized variables. As the original LoRA formulation is widely used in practice, these analyses do not address realistic implementations.
@@
Third, several works prove the convergence of LoRA under a strong boundedness assumption that artificially enforces Lipschitz smoothness. This includes federated LoRA algorithms [sun2024improving, park2025communicationefficientfederatedlowrankupdate, ghiasvand2025decentralizedlowrankfinetuninglarge], for which single-device LoRA is a special case. These works assume that the adapter matrix norms are uniformly bounded by constants. In addition, [jiang2024a] analyzes convergence of LoRA and representation fine-tuning (ReFT) within the same framework, assuming that the singular values of $A$ and $B$ are uniformly upper bounded, which upper bounds the matrix norms by rank-dependent constants that appear in the final convergence result. They also require that the gradient is uniformly bounded. These assumptions force Lipschitz smoothness in terms of the constants and are overly restrictive.
@@ 3.1 Preliminaries
where $\ell:\mathbb{R}^{m\times n}\to\mathbb{R}$ represents the loss function, $W_{0}\in\mathbb{R}^{m\times n}$ represents a frozen (pretrained) weight matrix, and $W\in\mathbb{R}^{m\times n}$ (sometimes denoted as $\Delta W$) represents the update to the frozen weights after fine-tuning. In the LoRA algorithm, we parametrize $W=BA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$ are the low-rank adapter matrices with rank $r<\min\{m,n\}$. During finetuning, only $A$ and $B$ are optimized while $W_{0}$ remains fixed. To simplify notation, we can reformulate (1) with the function $\mathcal{L}(W)=\ell(W_{0}+W)$ to obtain the LoRA minimization problem,
@@
$A_{t+1}=A_{t}-\eta_{t}\nabla_{A}\mathcal{L}(B_{t}A_{t}),\qquad B_{t+1}=B_{t}-\eta_{t}\nabla_{B}\mathcal{L}(B_{t}A_{t}).$ (3)
We require the following standard assumptions for our analysis.
Assumption 1 (Lipschitz smoothness).
@@
Equivalently, $\mathcal{L}$ satisfies the following “descent lemma”
@@
Assumptions 1 and 2 are fundamental to the study of gradient descent convergence. Critically, even with these assumptions, $\mathcal{L}$ is not Lipschitz smooth in $B$ or AA, preventing the analysis of (3) via classic optimization techniques.
@@ 3.2 Results
In the following, we detail three steps for proving convergence of LoRA. First, we reformulate the problem (2) as optimization over a single variable $V$ that contains the stacked matrices $B$ and $A^{T}$. We can rewrite the loss function in terms of the outer product $VV^{T}$. Second, we derive a modified “descent lemma,” analogous to the descent lemma for Lipschitz smooth functions (5), that enables descent in one step as long as the step size is small enough with respect to the parameter norm $\lVert V\rVert$ and gradient norm. The last step of the proof is ensuring that the step size does not decrease too quickly, thereby guaranteeing convergence.
@@
To recover $BA$, we need to extract the top right block from $VV^{T}$. Let $I_{n}$ represent the $n\times n$ identity matrix. Let $E_{1}=\begin{bmatrix}I_{m}&0_{m\times n}\end{bmatrix}$, $E_{1}\in\mathbb{R}^{m\times(m+n)}$ represent the “extractor” matrix that extracts the top $m$ rows from a $(m+n)\times(m+n)$ matrix when applied from the left, and let $E_{2}\penalty 10000\ =\penalty 10000\ \begin{bmatrix}0_{n\times m}&I_{n}\end{bmatrix}^{T}$, $E_{2}\in\mathbb{R}^{(m+n)\times n}$ represent the matrix that extracts the right $n$ columns when applied from the right. Then we have that
@@
We next define a new function $\mathcal{J}\penalty 10000\ :\penalty 10000\ \mathbb{R}^{(m+n)\times r}\to\mathbb{R}$ such that
@@
Step 2: “Lipschitz-like” descent lemma.
@@
The function $\mathcal{J}$ is not Lipschitz smooth, due to the multiplicative factor of $V$ in $\nabla\mathcal{J}(V)$. However, we can still obtain a “descent lemma” (Lemma 1) that is analogous to the classic descent lemma of Lipschitz smooth functions (5), but with more higher-order terms.
Lemma 1.
@@
$\mathcal{J}(V_{2})\leq$ $\mathcal{J}(V_{1})+\langle\nabla\mathcal{J}(V_{1}),V_{2}-V_{1}\rangle_{F}+\frac{2\sqrt{2}}{3}L\lVert V_{2}-V_{1}\rVert^{3}\lVert V_{1}\rVert+\sqrt{2}L\lVert V_{2}-V_{1}\rVert^{2}\lVert V_{1}\rVert^{2}+\frac{\sqrt{2}L}{3}\lVert V_{2}-V_{1}\rVert^{3}$
$+\frac{\sqrt{2}L}{4}\lVert V_{2}-V_{1}\rVert^{4}+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert V_{2}-V_{1}\rVert^{2}.$
@@
Step 3: Control step size to obtain convergence.
Based on Lemma 1, we can pick $\eta_{t}$ small enough to minimize the higher-order terms, thereby guaranteeing descent in one step as stated next.
Lemma 2.
@@
Lemma 2 states that to update $V_{t+1}$, we require the norm of the previous iterate $\lVert V_{t}\rVert$ and the gradient of the original loss function $\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$. By summing (9) over all t=0,,T1t=0,...,T-1 and telescoping sums, we determine that the minimum gradient norm over $t$ is upper bounded as
@@
However, (10) is not sufficient to conclude convergence; we also need to show that the sum $\sum_{t=0}^{T-1}\eta_{t}$ diverges as $T$ goes to infinity rather than converging to a finite value. Based on (8), this automatically holds if $\lVert V_{t}\rVert^{2}$ is uniformly upper bounded, leading to $\sum_{t=0}^{T-1}\eta_{t}=\Theta(T)$. However, if the parameter norm is not bounded, then $\lVert V_{t}\rVert^{2}$ may grow to infinity. In the proof of Theorem 3, we show that even in the worst case, we have $\lVert V_{t}\rVert^{2}=O(t)$. The sum $\sum_{t=0}^{T-1}\eta_{t}$ is therefore lower bounded by a harmonic series which is $\Theta(\log(T))$, leading to Theorem 3.
Theorem 3.
Suppose Assumption 1 and 2 hold. We perform $T$ steps of LoRA gradient descent (3), where at time step $t$ the step size $\eta_{t}$ is determined as

### v2

@@ Abstract
The low-rank adaptation (LoRA) algorithm for fine-tuning large models has grown popular in recent years due to its remarkable performance and low computational requirements. LoRA trains two “adapter” matrices that form a low-rank representation of the model parameters, thereby massively reducing the number of parameters that need to be updated at every step. Although LoRA is simple, its convergence is poorly understood due to the lack of Lipschitz smoothness, a key condition for classic convergence analyses. As a result, current theoretical results only consider asymptotic behavior or assume strong boundedness conditions which artificially enforce Lipschitz smoothness. In this work, we provide for the first time a non-asymptotic convergence analysis of the original LoRA gradient descent algorithm, which reflects widespread practice, without such assumptions. Our work relies on three key steps: i) reformulating the problem in terms of the outer product of the stacked adapter matrices, ii) a modified descent lemma for the “Lipschitz-like” reparametrized function, and iii) controlling the learning rate. With this approach, we prove that LoRA gradient descent converges to a stationary point at rate $O(\frac{1}{\log T})$, where $T$ is the number of iterations. We conduct numerical experiments to validate our theoretical findings.
Machine Learning, ICML
@@ 1 Introduction
Modern applications of large language models (LLMs) typically involve self-supervised pretraining of a large foundation model, which is later fine-tuned on a smaller task-specific dataset through supervised learning. Parameter efficient fine-tuning methods, which only train a small number of parameters, can efficiently adapt LLMs to a multitude of downstream tasks while maintaining comparable performance to fine-tuning of the full model parameters. One highly popular method is LoRA, or Low-Rank Adaptation (Hu et al., 2022), which trains a low-rank representation of the model parameters. Specifically, once the pretraining stage yields a model weight matrix $W_{0}$, LoRA only updates low-rank adapter matrices $A$, $B$, such that the final weights are $W_{0}+BA$. The matrices $A$ and $B$ can be updated with any optimization method, including gradient descent. However, even this simple algorithm is challenging to analyze; as observed in (Sun et al., 2024; Malinovsky et al., 2024), even if the original loss function is Lipschitz smooth, the new loss function— reparametrized in terms of $A$ and BB— is not. This nonsmoothness is a major obstacle to applying classic convergence analysis techniques to LoRA.
As a result, current theoretical analyses of LoRA fall into one of three categories. First, there are several works that examine LoRA in infinite regimes, such as its asymptotic convergence or its behavior on infinitely wide neural networks (Kim et al., 2025; Jang et al., 2024; Zeng and Lee, 2024; Yaras et al., 2024; Zhang and Pilanci, 2024; Hayou et al., 2024). However, these works do not establish a non-asymptotic convergence rate for finite models. Second, many works propose and analyze memory-efficient algorithms that resemble but do not address the original LoRA algorithm. These include GaLore (Zhao et al., 2024) and LoRA variants that only update a single adapter matrix at a time (Malinovsky et al., 2024; Sokolov et al., 2025). Third, some recent works analyze algorithmic frameworks that do extend to LoRA, and do derive a convergence rate (Jiang et al., 2024; Sun et al., 2024; Ghiasvand et al., 2025). However, these works require that the adapter matrices $A$ and $B$ are uniformly upper bounded by some constants which appear in the final convergence bound. These restrictive assumptions artificially enforce Lipschitz smoothness within the new parametrization, essentially reducing the problem to standard gradient descent with no substantially new proof techniques. Considering the existing literature, we seek to answer the question:
@@
In this work, we close the above research gap by establishing the first non-asymptotic convergence analysis of LoRA gradient descent, without requiring asynchronous updates, bounded adapter matrices, or Lipschitz smoothness of the reparametrized loss function. We only assume that the original loss function is Lipschitz smooth and lower bounded. Our novel proof approach involves stacking $B$ and $A$ into a single matrix $V$ and deriving a modified “Lipschitz-like” descent lemma for gradient descent with respect to $V$. We prove that to achieve descent at each step, the algorithm requires a learning rate that is “small enough” with respect to the norm of the current parameters and gradient. This recursive relationship between the parameters, gradient, and learning rate introduces instability in the convergence, resulting in an $O(\frac{1}{\log T})$ convergence rate to a stationary point. If we further assume that the adapter norms are bounded, we can recover the classic $O(\frac{1}{T})$ convergence rate of gradient descent on Lipschitz smooth functions.
By analyzing LoRA under minimal assumptions, we can make rigorous observations about its behavior. First, the convergence rate of LoRA does not (and should not) depend on the chosen rank, since gradient descent is a dimension-free algorithm. Second, we formalize the complex relationship between the parameter norm and the LoRA training dynamics, which has been alluded to in past work. We find that LoRA displays a “position dependency,” in that the convergence is slowed if the iterates are moving away from the origin and accelerated if they are moving towards the origin. This is due to the geometric changes introduced by the LoRA reparametrization, including the creation of a stationary point at the origin regardless of the structure of the original loss function.
Motivated by our theoretical insights, we perform experiments to examine the impact of learning rate choice under the LoRA reparametrization in practice. We train logistic regression and ResNet-18 models on the CIFAR-10 images dataset with adaptive and normalized learning rates that directly stem from theory, and we compare their performance against constant learning rates of similar values. We find that adjusting the learning rate based on the parameter norm or gradient norm can both accelerate and stabilize training by accounting for the unique structure induced by the low-rank reparametrization.
Our contributions can be summarized as follows:
  • We prove for the first time that LoRA gradient descent converges to a stationary point at rate $O(\frac{1}{\log T})$, only assuming that the original loss function is Lipschitz smooth and lower bounded.
  • Based on our theory, we propose and empirically validate practical methods for LoRA learning rate selection that demonstrate superior convergence.
In Section 2, we review the related work in detail and contextualize our contribution. In Section 3, we provide our convergence analysis, establishing preliminaries in Section 3.1, providing our main results in Section 3.2 and discussion in Section 3.3. Finally, in Section 4, we provide the results of our experiments. Code is open-sourced at the anonymous Github repository https://anonymous.4open.science/r/lora-C709/.
@@ 2 Related Work
The LoRA algorithm, first proposed in (Hu et al., 2022), is motivated by the idea of a viable “low-rank representation” of deep neural networks (Oymak et al., 2019; Li et al., 2018). In particular, (Aghajanyan et al., 2021) argues that large overparametrized models reside on a lower “intrinsic dimension,” and fine-tuning with a low dimension reparametrization, such as a random subset of the weight matrices, can yield strong performance while minimizing computational expenses.
@@
LoRA is simple, popular, and empirically effective for finetuning large language models. Many variants have been proposed to improve the initialization, performance, memory efficiency, or privacy of the algorithm (Shen, 2025; büyükakyüz2024oloraorthonormallowrankadaptation; Li et al., 2025; Meng et al., 2024; Wang et al., 2024), including but not limited to federated LoRA (Yang et al., 2025), quantized LoRA (QLoRA) (Dettmers et al., 2023), and ReLoRA (Lialin et al., 2024). While LoRA performs well in practice, its behavior in comparison to full-rank training is poorly understood, with empirical studies showing that the two approaches can produce completely different solutions (Shuttleworth et al., 2025; Biderman et al., 2024). Consequently, the theoretic properties of LoRA are highly relevant for understanding the strengths and weakness of this method.
@@
As highlighted in the introduction, the theory of LoRA convergence has been tackled from many angles. First, many works examine the behavior of LoRA in infinite regimes, whether by characterizing the kinds of solutions LoRA converges to at infinity or analyzing its convergence in the neural tangent kernel (NTK) regime, which defines the linearized training dynamics of an infinitely wide neural network (Malladi et al., 2023). For example, (Kim et al., 2025) argues that LoRA converges to a low-rank global minimum with high probability, assuming that one exists. The work (Zeng and Lee, 2024) studies the expressive power of the low-rank solutions achieved by LoRA. The work (Jang et al., 2024) shows that LoRA has no spurious local minima in the NTK regime. The LoRA+ (Hayou et al., 2024) algorithm, which updates the adapter matrices with differently scaled learning rates, is motivated by an NTK convergence analysis. Finally, the gradient flow dynamics of LoRA have been studied specifically for matrix factorization (Xu et al., 2025). These works do not establish a convergence rate for discrete-time LoRA for general functions.
@@
Second, there are memory-efficient LoRA-like algorithms with concrete convergence analyses that do not apply to the original LoRA formulation. These include GaLore, an algorithm which projects the gradient into a low-rank subspace but maintains a full-rank update of the parameters (Zhao et al., 2024), LDAdam (Robert et al., 2025), which performs adaptive optimization steps within lower dimensional subspaces, and Randomized Subspace Optimization (RSO) (Chen et al., 2025). The LoRA variant LoRA-One incorporates information from the full gradient at initialization (Zhang et al., 2025), and it is analyzed only for Gaussian data and mean-squared loss. Finally, the variants RAC-LoRA (Malinovsky et al., 2024) and Bernoulli-LoRA (Sokolov et al., 2025) only train one adapter matrix while freezing the other as a fixed projection matrix, essentially preserving Lipschitz smoothness while failing to capture the complex training dynamics of actual LoRA implementations. In practice, the LoRA matrices are updated simultaneously, causing nonlinearity and nonsmoothness in the optimized variables. As the original LoRA formulation is much more commonly used in practice, these analyses do not address realistic implementations.
@@
Third, several works prove the convergence of LoRA under a strong boundedness assumption that artificially enforces Lipschitz smoothness. This includes analyses of federated LoRA algorithms (Sun et al., 2024; Park and Klabjan, 2025; Ghiasvand et al., 2025), for which single-device LoRA is a special case. These works assume that the adapter matrix norms are uniformly bounded by constants. In addition, (Jiang et al., 2024) analyzes convergence of LoRA and representation fine-tuning (ReFT) within the same framework, assuming that the singular values of $A$ and $B$ are uniformly upper bounded, which upper bounds the matrix norms by rank-dependent constants that appear in the final convergence result. They also require that the gradient is uniformly bounded.
@@ 3.1 Preliminaries
where $\ell:\mathbb{R}^{m\times n}\to\mathbb{R}$ represents the loss function, $W_{0}~\in~\mathbb{R}^{m\times n}$ represents a frozen (pretrained) weight matrix, and $W\in\mathbb{R}^{m\times n}$ (sometimes denoted as $\Delta W$) represents the update to the frozen weights after fine-tuning. In the LoRA algorithm, we parametrize $W=BA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$ are the low-rank adapter matrices with rank $r<\min\{m,n\}$. During finetuning, only $A$ and $B$ are optimized while $W_{0}$ remains fixed. To simplify notation, we can reformulate (1) with the function $\mathcal{L}(W)=\ell(W_{0}+W)$ to obtain the LoRA minimization problem,
@@
$A_{t+1}=A_{t}-\eta_{t}\nabla_{A}\mathcal{L}(B_{t}A_{t}),$
$B_{t+1}=B_{t}-\eta_{t}\nabla_{B}\mathcal{L}(B_{t}A_{t}),$ (3)
where $\eta_{t}$ denotes the learning rate at time $t$. We require the following standard assumptions for our analysis.
Assumption 3.1 (Lipschitz smoothness).
@@
Equivalently, $\mathcal{L}$ satisfies the following descent lemma,
@@
Assumptions 3.1 and 3.2 are fundamental to the study of gradient descent convergence. Critically, even with these assumptions, $\mathcal{L}$ is not Lipschitz smooth in $B$ or AA, preventing the analysis of (3) via classic optimization techniques.
@@ 3.2 Results
In the following, we detail three steps for proving convergence of LoRA. First, we reformulate the problem (2) as optimization over a single variable $V$ that contains the stacked matrices $B$ and $A^{T}$. We can rewrite the loss function in terms of the outer product $VV^{T}$. Second, we derive a modified descent lemma, analogous to the descent lemma for Lipschitz smooth functions (5), that enables descent in one step as long as the learning rate is small enough with respect to the parameter norm $\lVert V\rVert$ and gradient norm. The last step of the proof is ensuring that the learning rate does not decrease too quickly, thereby guaranteeing convergence.
@@
To recover $BA$, we need to extract the top right block from $VV^{T}$. Let $I_{n}$ represent the $n\times n$ identity matrix. Let $E_{1}=\begin{bmatrix}I_{m}&0_{m\times n}\end{bmatrix}$, $E_{1}\in\mathbb{R}^{m\times(m+n)}$ represent the “extractor” matrix that extracts the top $m$ rows from a $(m+n)\times(m+n)$ matrix when applied from the left, and let $E_{2}~=~\begin{bmatrix}0_{n\times m}&I_{n}\end{bmatrix}^{T}$, $E_{2}\in\mathbb{R}^{(m+n)\times n}$ represent the matrix that extracts the right $n$ columns when applied from the right. Then we have that
@@
We next define the function $\mathcal{J}~:~\mathbb{R}^{(m+n)\times r}\to\mathbb{R}$ such that
@@
Step 2: Descent lemma.
@@
The function $\mathcal{J}$ is not Lipschitz smooth, due to the multiplicative factor of $V$ in $\nabla\mathcal{J}(V)$. However, we can still obtain a descent lemma (Lemma 3.3) that is analogous to the classic descent lemma of Lipschitz smooth functions (5), but with more higher-order terms.
Lemma 3.3.
@@
$\mathcal{J}(V_{2})\leq$ $\mathcal{J}(V_{1})+\langle\nabla\mathcal{J}(V_{1}),V_{2}-V_{1}\rangle_{F}$
$+\sqrt{2}L\lVert V_{2}-V_{1}\rVert^{2}\lVert V_{1}\rVert^{2}+\frac{\sqrt{2}L}{3}\lVert V_{2}-V_{1}\rVert^{3}$
$+\frac{2\sqrt{2}}{3}L\lVert V_{2}-V_{1}\rVert^{3}\lVert V_{1}\rVert+\frac{\sqrt{2}L}{4}\lVert V_{2}-V_{1}\rVert^{4}$
$+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert V_{2}-V_{1}\rVert^{2}.$
@@
Step 3: Control learning rate to achieve convergence.
Based on Lemma 3.3, we can pick $\eta_{t}$ small enough to minimize the higher-order terms, thereby guaranteeing descent in one step as stated next.
Lemma 3.4.
@@
Lemma 3.4 states that to update $V_{t+1}$, we require the norm of the previous iterate $\lVert V_{t}\rVert$ and the gradient of the original loss function $\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$. By summing (9) over all t=0,,T1t=0,...,T-1 and telescoping sums, we determine that the minimum gradient norm over $t$ is upper bounded as
@@
However, (10) is not sufficient to conclude convergence; we also need to show that the sum $\sum_{t=0}^{T-1}\eta_{t}$ diverges as $T$ goes to infinity rather than converging to a finite value. Based on (8), this automatically holds if $\lVert V_{t}\rVert^{2}$ is uniformly upper bounded, leading to $\sum_{t=0}^{T-1}\eta_{t}=\Theta(T)$. However, if the parameter norm is not bounded, then $\lVert V_{t}\rVert^{2}$ may grow to infinity. In the proof of Theorem 3.5, we show that even in the worst case, we have $\lVert V_{t}\rVert^{2}=O(t)$. The sum $\sum_{t=0}^{T-1}\eta_{t}$ is therefore lower bounded by a harmonic series which is $\Theta(\log(T))$, leading to Theorem 3.5.
Theorem 3.5.
Suppose Assumption 3.1 and 3.2 hold. We perform $T$ steps of LoRA gradient descent (3), where at time step $t$ the learning rate $\eta_{t}$ is determined as